What is the way to solve a basic equation? There are lots of steps, and I dont get them.
for example, what are the steps to solving the following equation:
2x + -10 = 10 + -3x
And can you explain the "why" in every step?

Answer by nerdybill(7384)

The idea is to isolate the unknown (x) to one side of the equation -- doesn't matter which side.
.
2x + -10 = 10 + -3x
Start by adding 10 to both sides:
2x + -10 +10 = 10 + -3x +10
simplifying:
2x = 20 + -3x
.
add 3x to both sides:
2x +3x = 20 + -3x +3x
simplifying:
5x = 20
.
finally, divide both sides by 5:
(5x)/5 = 20/5
simplifying:
x = 4
.
To check, plug the answer back into the original equation.